OpenStudy (ffasinger):

The slope of the line below is -2. Which of the following is the point-slope form of the line? A. y + 8 = -2(x - 4) B. y - 4 = -2(x + 8) C. y - 8 = -2(x + 4) D. y + 4 = -2(x - 8)

OpenStudy (nubeer):

use the formula \[y-y _{1}=m(x-x _{1})\] plut (4,-8) in this equation.. and value of m is also given to you

OpenStudy (abdullahm):

Hint: \(\bf\Large y-y_1 = m (x-x_1)\) m = slope = -2 \(\bf\Large (-4,8)\\\bf\Large (x_1,y_1)\)

OpenStudy (aishalicious1):

U still on here? 18 hours later?